find p and q if the vectors 3i+2pj-qk and 5i+6j+4k parallel
 
 

Grade:12th pass

1 Answers

Adarsh Shivam
2 years ago
As if two vectors are parallel, then their cross product will be zero:
(3i + 2pj - qk) × (5i +6j +4k) = 0
(8p + 6q)i - (12 + 5q)j + (18 - 10p)k = 0
8p + 6q = 0;
12 + 5q = 0;
18 - 10p= 0.
"p" = 9/5
"q" = -12/5